OpenStudy (anonymous):

Classifying 3x^5-8x^3-2x^2+5 by numbers of terms A. Trinomial B. Polynomial C. Binomial D. Polynomial of 5 terms

OpenStudy (anonymous):

I think D

OpenStudy (anonymous):

Your expression has got 4 terms. That makes it a polynomial. B is the answer.

OpenStudy (anonymous):

@Hoslos Thanks, wasn't sure. : (

OpenStudy (anonymous):

Thanx! :) and it's ok cx

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Consider the leading term of the polynomial function. What is the end behavior of the graph? 4x^5+1x

OpenStudy (anonymous):

I'll add the choices in a sec

OpenStudy (anonymous):

Time up if you are counting.

OpenStudy (anonymous):

A. Since n is odd and a is positive, the end behavior is up and down B. down and up C. Down and down D. Up and up

OpenStudy (anonymous):

up and up then?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Okie :)

OpenStudy (anonymous):

It keeps going up.

OpenStudy (anonymous):

Use synthetic division to find P(3) for P(x)=x^4-6x^3-4x^2-6x-2 A. 3 B. -137 299 -47

OpenStudy (anonymous):

B

OpenStudy (anonymous):

How many more? Just curious.

OpenStudy (anonymous):

As many as you're willing to do ;-; I have finals tomorrow morning at 8 am and it's 3 am already and I have 5 more assignments to do Dx

OpenStudy (anonymous):

WOW O-O

OpenStudy (anonymous):

I know Dx and I have work and ugh. The struggle

OpenStudy (anonymous):

Ok. Next question!

OpenStudy (anonymous):

Alrighty ^^ one sec

OpenStudy (anonymous):

Write the polynomial in factored form. x^3 – 3x^2 – 10x

OpenStudy (anonymous):

13. What are the real or imaginary solutions of the polynomial equation? 27x^3 + 125 = 0

OpenStudy (anonymous):

First one: x(x+2)(x-5) Second one: I don't really know try over here: https://mathway.com/ my interneti= is slowing down a bit and won't load. Sorry...

OpenStudy (anonymous):

It's ok, I think I got one more then I'll be done for now, I'm sorry :(

OpenStudy (anonymous):

\[x^3=-125/27rightarrowx=\sqrt[3]{-125/27}=-5/3\] A real number.

OpenStudy (anonymous):

Thank Hoslos.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

@Squeeks1713 It's ok! Throw the questions!

OpenStudy (anonymous):

It is okay.

OpenStudy (anonymous):

Use the Remainder Theorem to find the remainder when P(x) = x^4 – 9x^3 – 5x^2 – 3x + 4 is divided by x + 3.

OpenStudy (anonymous):

I think the answer is 292.

OpenStudy (anonymous):

To solve this, you have multiply one term that when multiplied by x from x+3 gives x^4. Then you make subtraction of like terms. One second step, multiply a number with x from x+3 that gives -9x^3 and so on.

OpenStudy (anonymous):

So confusing lol I don't know how you guys do this omg

OpenStudy (anonymous):

I just used mathway!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Bye

OpenStudy (anonymous):

Let me know if you need help.

OpenStudy (anonymous):

I will, thank you!

OpenStudy (anonymous):

Bye

OpenStudy (anonymous):

The answer is 292 using the aforementioned method. @Lovelarap got it as well.

OpenStudy (anonymous):

Cool!